OpenStudy (anonymous):

A = P(1 + rt). Find t if A = 168, P = 150, and r = 0.08.

OpenStudy (anonymous):

1.5

OpenStudy (anonymous):

just substitute 168=150(1+0.08t) divide by 150 =168/150=1.12 1.12=1+0.08t 1.12-1=0.08t 0.12=0.08t divide by 0.08 t=0.12/0.08 t=1.5